OpenStudy (anonymous):

Write the equation in slope form and identify the slope and y-intercept. 3x+2y=17 The equation in slope-intercept form is y= The slope is? The y-intercept is?

hero (hero):

Hint: Isolate y

OpenStudy (anonymous):

I need to know how to get the answer. I need the steps. I have 3 more of these two do.

hero (hero):

1. Subtract 3x from both sides. Let me know what you get.

OpenStudy (anonymous):

Ok hold on a second

OpenStudy (anonymous):

14

hero (hero):

You subtracted 3 from the right side, but I wanted you to subtract 3x from both sides and let me know what the equation would look like afterwards

OpenStudy (anonymous):

2y=14

OpenStudy (anonymous):

7y

hero (hero):

If you subtracted 3x from both sides you would get this 3x - 3x + 2y = 17 - 3x Then you would simplify to get this: 2y = 17 - 3x

hero (hero):

The next step is to divide both sides by 2. Let me know what you get.

OpenStudy (anonymous):

y=7

OpenStudy (anonymous):

Is this right

hero (hero):

Not exactly

OpenStudy (anonymous):

What did I do wrong

OpenStudy (anonymous):

Could u show me an example to go by

hero (hero):

You were supposed to divide both sides by 2. Instead, you posted y = 7.

hero (hero):

Where's your math book?

OpenStudy (anonymous):

I don't have one its through online

hero (hero):

You should probably go get one. I have one I can give you.

OpenStudy (anonymous):

could you show me how to do this one so I will have example to go by to do the other ones I got to solve.

hero (hero):

Yes, I will give you a book that has plenty examples.

OpenStudy (anonymous):

Thats not going to do me no good right know. It will take days for a book to get here.

hero (hero):

No it won't. I will give you a book that you can download.